XYZ has 100 000 common shares and issues 50 000 preferred shares in their first round of financing at R10, which is the price at which common shares are trading in the market currently. In a subsequent financing round the company issues new shares at R7.50. The adjusted conversion price, using the weighted-average anti-dilution provision formula, is closest to:
Answer #1

Adjusted conversion price C2 = C1 x (A + B) / (A + C)

Where C2 = After conversion price

C1 = Before conversion price = 10

A = number of shares before a new issue = 100000

B = total consideration based on C1 = 750000 /10 = 75000

Where 750000 come from =100000(Current Shares)*7.5 (New Price)

C = new shares issued = 100000

= 10* (100000+75000) / (100000+100000) = 8.75

So the answer is 8.75